According to the record of the registrar’s office at a state university, 35% of the students are freshman, 25% are sophomore, 16% are junior and the rest are senior. Among the freshmen, sophomores, juniors and seniors, the portion of students who live in the dormitory are, respectively, 80%, 60%, 30% and 20%


Referring to the information above, what is the probability that a randomly selected student is a sophomore who does not live in a dormitory?
Answer #1

P(sophomore who does not live in a dormitory ) = P(sophomore)*P(does not live in dormitory ) = 0.25*(1-0.60)=0.10